OpenStudy (anonymous):

In the following problems, determine if the series converges. If it does, is the convergence conditional or absolute? 1) \[\sum_{n=1}^{\infty} e ^{1/n}/n ^{2} \] a)using alternating test b)using integral test 2) \[\sum_{n=1}^{\infty} (−1)n \cos(\pi n) \] 3) \[\sum_{n=1}^{\infty} (-1)^{n}/n!\] 4) \[\sum_{n=1}^{\infty}2^{n}n!/n ^{n} \]

OpenStudy (anonymous):

1) \[\sum_{n=1}^{\infty} e ^{1/n}/n ^{2} \] a) using alternating test b) using integral test 2) \[\sum_{n=1}^{\infty} (−1)n \cos(\pi n) \] 3) \[\sum_{n=1}^{\infty} (-1)^{n}/n!\] 4) \[\sum_{n=1}^{\infty}2^{n}n!/n ^{n} \]